OpenStudy (anonymous):

In (triangle ) ABC, AB = x, BC = y, and CA = 2x. A similarity transformation with a scale factor of 0.5 maps (triangle) ABC to (triangle) MNO, such that vertices M, N, and O correspond to A, B, and C, respectively. If OM = 5, what is AB?

OpenStudy (anonymous):

@Eli_Moses

OpenStudy (eli_moses):



OpenStudy (eli_moses):



OpenStudy (anonymous):

10

OpenStudy (eli_moses):

no remember that the scale factor is 0.5 so figure 1 is dilated by 0.5 to create figure 2

OpenStudy (anonymous):

so it d be 5

OpenStudy (eli_moses):

yea try it out before you put it

OpenStudy (anonymous):

ok so it is 5 correct

OpenStudy (anonymous):

@Eli_Moses

OpenStudy (eli_moses):

yea

jimthompson5910 (jim_thompson5910):

AC maps to MO. Since we're applying a scale factor of 0.5, this means MO = 0.5*AC MO = 0.5*2x 5 = 0.5*2x 5 = x x = 5 So that's why AB = 5

OpenStudy (anonymous):

ok great